# State-of-the-Art Approaches to Enhancing Privacy Preservation of Machine Learning Datasets: A Survey

### v1

@@ 1 Introduction
Recent developments in the field of machine learning, especially in the domain of deep learning, have markedly impacted numerous sectors, such as next-g network[69], financial technology, and surveillance systems[13, 12], signaling a transformative phase in these industries. Simultaneously, the surge in machine learning-powered artificial intelligence has brought privacy issues to the forefront. Hence, the notion of Privacy-preserving Machine Learning (PPML) has arisen as a prominent area of interest in both the industrial and academic communities.
@@
In the field of PPML, a primary focus is on ensuring that the implemented ML models prevent the escape of confidential information from the training data beyond the trusted boundaries of the data sources. Specifically, during training phase, the main issues of privacy leakage are centered around the handling of data and its computation. Current research addresses these challenges through two key approaches: (i) determining methods to refine or filter the training data with the aim of either minimizing or entirely eradicating any information sensitive to privacy; and (ii) developing techniques for processing the training data in a way that upholds privacy.
This report concisely examines the attacks associated with privacy leakage of training data in machine learning, and delve into the techniques employed for preserving training data privacy in both centralized Learning and collaborative learning to counter these attacks. Through an analysis of these techniques, the report seeks to offer an in-depth understanding of effective strategies for safeguarding training data in machine learning against privacy violations. Additionally, it aims to elucidate the balance between privacy and utility or efficiency that arises in the context of these privacy-preserving approaches.
Privacy-Preserving Techniques. Privacy-preserving models in machine learning are fundamentally designed to not disclose extra information about individuals included in their training data. This aspect is of paramount importance for ML systems based on sensitive data, including digital assets[31] and medical image processing applications  [27]. The protection of training data in ML involves a strategic interplay between cryptographic techniques and Differential Privacy (DP). Besides, few works utilizing Trusted Execution Environment (TEE) or SGX to achieve PPML  [23].
@@ 3 Privacy-Preserving Techniques in PPML
Regarding training data privacy in privacy-preserving machine learning , current approaches primarily focus on several strategies:
@@ 3.2.1 Model Training
In response to this challenge, an improved version named ABY2.0 [48] was proposed. ABY2.0 enhanced the performance of the original ABY framework by shifting the bulk of OTs to the preprocessing phase. This meant that the time-consuming process of executing OTs did not hinder the efficiency of the actual computation phase. Furthermore, ABY2.0 supported multi-input multiplication operations without increasing the online communication overhead, thereby making it more efficient for complex computations.
@@
Adopting heterogeneous computing architectures can make defending against adversarial attacks in complex systems more viable and cost-effective[31]. By integrating various computing resources such as CPUs, GPUs, and FPGAs, security protocols and cryptographic algorithms can be expedited, improving defense strategies’ efficiency and practicality[74]. This method optimizes the allocation of computational tasks to the most appropriate processors, enhancing both performance and energy efficiency[17, 70]. Consequently, heterogeneous computing is crucial for implementing advanced, affordable security measures in safeguarding against adversarial threats[72, 30].
@@ 3.3 Hardware-protected Method
Trusted Execution Environments (TEEs) like Intel SGX are engineered to enable the secure execution of programs within hardware-protected enclaves. Intel SGX specifically forms secure memory regions within the address space, segregating these areas from all other system codes. This technique facilitates both authentication and encapsulation, allowing external entities to verify that communications originate from an authentic enclave. Within the framework of Privacy-Preserving Machine Learning (PPML) protocols that utilize TEEs, there is an essential process whereby involved parties must transfer confidential data or models into an enclave through a protected channel. Following the authentication of the software operating within the enclave, it can securely process and return encrypted results. These TEE-based methods are relevant and useful in various aspects of machine learning, encompassing both the training phase and the prediction stage of models, as indicated in [22, 23, 43]. The use of TEEs in PPML ensures that the computation of sensitive data or models is performed in a secure and isolated environment, thus enhancing privacy and security.
@@ 4 Defenses in Collaborative/Federated Learning
While Federated Learning (FL) contributes to privacy, it does not entirely prevent data exposure. Studies have shown that inversion attacks might reveal detailed images from models’ weights or gradients [18, 62, 61, 73]. Additionally, when models are involved in MLaaS, they are vulnerable to misuse or theft by untrustworthy entities. To overcome these shortcomings, FL needs to be enhanced with additional privacy-preserving methods. Implementing FL with secure aggregation (SecAgg) or integrating differential privacy (DP) can impede attacks aimed at dataset reconstruction. Furthermore, applying secure multi-party computation (SMPC) protocols can safeguard the models during the inference process.
@@ 5 Comparison and Summary
    1. (a)
      Performance: Network latency can affect SMPC, and its efficiency is also contingent on the specific protocols or computations being executed.
    2. (b)
      Privacy: Trust among participants is essential for SMPC to function correctly.
    3. (c)
      Application: It is typically used in collaborative or distributed learning environments where multiple parties compute a function without revealing their inputs.
@@
    1. (a)
      Performance: Zero-Knowledge Proofs aim for smaller proofs to enhance efficiency, and some protocols support parallel computation.
    2. (b)
      Privacy: The level of privacy varies with the chosen protocol.
    3. (c)
      Application: ZKPs are often used for data validation processes to prove data integrity without revealing the data itself.
@@
    1. (a)
      Performance: HE can be computationally intensive, and the performance depends on the chosen scheme and parameters.
    2. (b)
      Privacy: It offers provable security by allowing computations on encrypted data.
    3. (c)
      Application: HE is ideal for outsourced or centralized computing scenarios, enabling computation on encrypted data stored in the cloud.
@@
    1. (a)
      Performance: The overhead is significantly impacted by the processes of key generation and management.
    2. (b)
      Privacy: Data access can be fine-grained, but there may be some inevitable leakage of information.
    3. (c)
      Application: It’s used in situations requiring fine-grained access control over encrypted data.
@@
    1. (a)
      Performance: The approach is accuracy-limited due to the introduction of noise.
    2. (b)
      Privacy: Privacy guarantees are noise-dependent, balancing data utility with privacy.
    3. (c)
      Application: Commonly applied in data analysis and aggregation to protect individual data points within a dataset.
@@
    1. (a)
      Performance: Overhead is influenced by the setup and management of secure enclaves.
    2. (b)
      Privacy: Ensures isolation of code and data but may be susceptible to side-channel attacks.
    3. (c)
      Application: Used within hardware environments to secure sensitive computations.

### v2

@@ 1 Introduction
Recent developments in the field of machine learning, especially in the domain of deep learning, have markedly impacted numerous sectors, such as next-g network[77, 85, 86], financial technology[34], marketing[63, 64], and surveillance systems[14, 13, 78], signaling a transformative phase in these industries. Simultaneously, the surge in machine learning-powered artificial intelligence has brought privacy issues to the forefront. Hence, the notion of Privacy-preserving Machine Learning (PPML) has arisen as a prominent area of interest in both the industrial and academic communities.
@@
In the field of PPML, a primary focus is on ensuring that the implemented ML models prevent the escape of confidential information from the training data beyond the trusted boundaries of the data sources. Specifically, during the training phase, the main issues of privacy leakage are centered around the handling of data and its computation. Current research addresses these challenges through two key approaches: (i) determining methods to refine or filter the training data with the aim of either minimizing or entirely eradicating any information sensitive to privacy; and (ii) developing techniques for processing the training data in a way that upholds privacy.
This report concisely examines the attacks associated with privacy leakage of training data in machine learning and delves into the techniques employed for preserving training data privacy in both centralized Learning and collaborative learning to counter these attacks. Through an analysis of these techniques, the report seeks to offer an in-depth understanding of effective strategies for safeguarding training data in machine learning against privacy violations. Additionally, it aims to elucidate the balance between privacy and utility or efficiency that arises in the context of these privacy-preserving approaches.
Privacy-Preserving Techniques. Privacy-preserving models in machine learning are fundamentally designed to not disclose extra information about individuals included in their training data. This aspect is of paramount importance for ML systems based on sensitive data, including digital assets[34] and medical image processing applications  [28]. The protection of training data in ML involves a strategic interplay between cryptographic techniques and Differential Privacy (DP). Besides, few works utilize a Trusted Execution Environment (TEE) or SGX to achieve PPML  [24].
@@ 3 Privacy-Preserving Techniques in PPML
Regarding training data privacy in privacy-preserving machine learning , current approaches [31] primarily focus on several strategies:
@@ 3.2.1 Model Training
bIn response to this challenge, an improved version named ABY2.0 [51] was proposed. ABY2.0 enhanced the performance of the original ABY framework by shifting the bulk of OTs to the preprocessing phase. This meant that the time-consuming process of executing OTs did not hinder the efficiency of the actual computation phase. Furthermore, ABY2.0 supported multi-input multiplication operations without increasing the online communication overhead, thereby making it more efficient for complex computations.
@@
Adopting heterogeneous computing architectures can make defending against adversarial attacks in complex systems more viable and cost-effective[2, 32, 34, 70, 69, 86]. By integrating various computing resources such as CPUs, GPUs, and FPGAs, security protocols and cryptographic algorithms can be expedited, improving defense strategies’ efficiency and practicality[83]. This method optimizes the allocation of computational tasks to the most appropriate processors, enhancing both performance and energy efficiency[18, 79]. Consequently, heterogeneous computing is crucial for implementing advanced, affordable security measures in safeguarding against adversarial threats[zhang2019optimization, 33].
@@ 3.3 Hardware-protected Method
Trusted Execution Environments (TEEs) like Intel SGX are engineered to enable the secure execution of programs within hardware-protected enclaves. Intel SGX specifically forms secure memory regions within the address space, segregating these areas from all other system codes. This technique facilitates both authentication and encapsulation, allowing external entities to verify that communications originate from an authentic enclave. Within the framework of Privacy-Preserving Machine Learning (PPML) protocols that utilize TEEs, there is an essential process whereby involved parties must transfer confidential data or models into an enclave through a protected channel. Following the authentication of the software operating within the enclave, it can securely process and return encrypted results. These TEE-based methods are relevant and useful in various aspects of machine learning, encompassing both the training phase and the prediction stage of models, as indicated in [23, 24, 46]. The use of TEEs in PPML ensures that the computation of sensitive data or models is performed in a secure and isolated environment, thus enhancing privacy and security [68].
@@ 4 Defenses in Collaborative/Federated Learning
While Federated Learning (FL) contributes to privacy, it does not entirely prevent data exposure. Studies have shown that inversion attacks might reveal detailed images from models’ weights or gradients [19, 67, 66, 82, 81]. Additionally, when models are involved in MLaaS, they are vulnerable to misuse or theft by untrustworthy entities. To overcome these shortcomings, FL needs to be enhanced with additional privacy-preserving methods. Implementing FL with secure aggregation (SecAgg) or integrating differential privacy (DP) can impede attacks aimed at dataset reconstruction. Furthermore, applying secure multi-party computation (SMPC) protocols can safeguard the models during the inference process.
@@ 5 Comparison and Summary
    (a) Performance: Network latency can affect SMPC, and its efficiency is also contingent on the specific protocols or computations being executed.
    (b) Privacy: Trust among participants is essential for SMPC to function correctly.
    (c) Application: It is typically used in collaborative or distributed learning environments where multiple parties compute a function without revealing their inputs.
@@
    (a) Performance: Zero-Knowledge Proofs aim for smaller proofs to enhance efficiency, and some protocols support parallel computation.
    (b) Privacy: The level of privacy varies with the chosen protocol.
    (c) Application: ZKPs are often used for data validation processes to prove data integrity without revealing the data itself.
@@
    (a) Performance: HE can be computationally intensive, and the performance depends on the chosen scheme and parameters.
    (b) Privacy: It offers provable security by allowing computations on encrypted data.
    (c) Application: HE is ideal for outsourced or centralized computing scenarios, enabling computation on encrypted data stored in the cloud.
@@
    (a) Performance: The overhead is significantly impacted by the processes of key generation and management.
    (b) Privacy: Data access can be fine-grained, but there may be some inevitable leakage of information.
    (c) Application: It’s used in situations requiring fine-grained access control over encrypted data.
@@
    (a) Performance: The approach is accuracy-limited due to the introduction of noise.
    (b) Privacy: Privacy guarantees are noise-dependent, balancing data utility with privacy.
    (c) Application: Commonly applied in data analysis and aggregation to protect individual data points within a dataset.
@@
    (a) Performance: Overhead is influenced by the setup and management of secure enclaves.
    (b) Privacy: Ensures isolation of code and data but may be susceptible to side-channel attacks.
    () Application: Used within hardware environments to secure sensitive computations.